### v2

# Ordinary Least Squares as an Attention Mechanism

Abstract
I show that ordinary least squares (OLS) predictions can be rewritten as the output of a restricted attention module, akin to those forming the backbone of large language models. This connection offers an alternative perspective on attention beyond the conventional information retrieval framework, making it more accessible to researchers and analysts with a background in traditional statistics. It falls into place when OLS is framed as a similarity-based method in a transformed regressor space, distinct from the standard view based on partial correlations. In fact, the OLS solution can be recast as the outcome of an alternative problem: minimizing squared prediction errors by optimizing the embedding space in which training and test vectors are compared via inner products. Rather than estimating coefficients directly, we equivalently learn optimal encoding and decoding operations for predictors. From this vantage point, OLS maps naturally onto the query-key-value structure of attention mechanisms. Building on this foundation, I discuss key elements of Transformer-style attention and draw connections to classic ideas from time series econometrics.
1 Introduction
The self-attention mechanism popularized by vaswani2017attention has sparked a technical revolution that fundamentally reshaped the landscape of deep learning and propelled AI into the public eye. Its introduction in 2017 single-handedly ushered in the era of large language models (LLMs), shifting the natural language processing field away from convolutional/recurrent neural network architectures toward Transformer-based models.
Yet, for those with a background in applied statistics and machine learning, the formulation of attention in terms of queries, keys, and values—borrowed from information retrieval systems—may be elusive. In this paper, I show that test-set predictions from a linear regression estimated via ordinary least squares (OLS) can be rewritten as the output of a simplified attention module, where the weight matrices admit a closed-form solution. This connection offers an alternative understanding of OLS and a perspective on attention that will resonate with analysts familiar with regression methods.
The Double Life of Least Squares. The first arc of the paper introduces an alternative view of OLS. Standard OLS predictions are usually expressed as combinations of coefficients and regressors values. Yet, another mathematically equivalent interpretation exists: OLS is also a similarity-based estimator in a transformed feature space. Predictions are linear combinations of training observations of the target variable, with weights determined by inner products between test and training vectors encoded as mutually orthogonal factors. Observations closer to the target receive greater weight or, said differently, more attention. Hence, OLS is as much correlation-based as it is proximity-based.
We can go further. OLS estimation itself can be viewed as optimizing the encoding and decoding of input vectors rather than estimating coefficients. That is, OLS minimizes squared prediction errors by determining the optimal embedding space in which training and test vectors are compared by simple inner products. For OLS, this optimal embedding is available in closed-form and corresponds to the inverse covariance matrix of predictors. From this vantage point, mapping OLS into the information retrieval framework is straightforward: queries are encoded out-of-sample predictors, keys are encoded in-sample predictors, values correspond to realized dependent variable outcomes, and the attention weighting function reduces to the identity.
Under the Hood of Transformers. The second arc builds on the established correspondence. I revisit core elements of the attention mechanism in Transformer architectures, including dimensionality reduction, nonlinearity, shrinkage, and multi-head attention. I then explore how these components interact within the Transformer sequence, drawing links with time series econometrics. Specifically, I connect self-attention and vector autoregressions by highlighting their mutual reliance on conditional expectations as filters, eliminating the need for sequential estimation and ensuring a model-consistent notion of context. This analogy opens a gateway to clarify the roles of word embeddings, training over millions of sequences, and the recursive chaining of attention and feed-forward modules. Finally, I discuss masking—a method employed in language models to prevent information leakage from future to past—and relate it explicitly to pseudo-out-of-sample evaluation.
Related Work. The connection between attention mechanisms—based on a nonlinearly transformed dot product in a latent space—and kernel methods has been widely discussed in the literature (tsai2019transformer; choromanski2020rethinking; peng2021random). However, the parallel is often limited to a similarity-based interpretation of attention weights, without fully crossing the bridge to a correlation-based perspective. In this paper, I show that this gap can be bridged by leveraging an alternative embedding-based view of OLS predictions.
From the OLS predictions perspective, perhaps the closest formulation to the one presented in this paper is that of the representer theorem, which states that solutions to nonparametric ridge regression can be expressed as a finite combination of kernelized inner products between the new data vector and training samples (scholkopf2001generalized; wahba2019representer). This link is further discussed in dual. However, these formulations always feature a vector of observation-specific weights which serves as a set of parameters defining the dual solution to regularized least squares problems. The precise interpretation of the parameters’ role, beyond serving as linear weights for kernelized inner products in the dual representation of predictions, remains unclear.
In this paper’s formulation, where OLS is reframed as an attention module, weights on training observations emerge directly from comparing training and test data in a transformed, orthonormal space. This transformation enables the direct assignment of weights to training outcomes, analogous to “values” in the information retrieval framework.
Two recent studies have drawn connections between attention mechanisms and regression tasks. marion2024attention introduced the single-location regression task, a simplified setting where only one token in a sequence determines the output, with its position treated as a latent variable. They demonstrated that a non-linear self-attention layer can effectively identify the relevant token and perform the particular regression. garnelo2023exploring expanded on this by examining key-value-query models, which share the input structure of attention mechanisms but differ in computation. They introduced a new module that generalizes linear attention and computes regularized least squares solutions within the key-value-query framework. The present paper focuses on the standard attention mechanism and explains why similarities between regression solutions and attention modules arise: least squares-based predictions are also interpretable as inner-product-based similarity computations in an optimized orthonormal space–an interpretation closely aligned with how keys, queries, and values interacts in LLMs.
kwon2024large, ludwig2025large, and others have taken an econometric perspective on LLMs, studying how outputs respond to prompt strategies and outlining best practices for empirical research using them. This paper has a different focus. Rather than treating LLMs as given black-box tools, the objective is to provide a statistical interpretation of the mechanisms operating behind the scenes. Recent contributions in this direction include kelly2025 and hasan, who study attention in the context of portfolio construction and provide sharper insights for the linear case. However, this paper is the first to point out and exploit the equivalence between traditional regression methods and (restricted) attention mechanisms.
Outline. The rest of this paper is organized as follows. Section 2 presents an alternative view on OLS, reviews the attention mechanism, and connects the two under some conditions. Section 3 discusses extensions of these ideas to incorporate many other features of the actual attention mechanism employed in LLMs. Section 4 leverages the nonlinear regression interpretation of attention to unpack key elements of the Transformer architecture beyond the attention module itself, drawing connections to time series econometrics. Section 5 provides experimental evidence on the predictive performance of the nonlinear Attention Regression framework across a range of data-generating processes. Section 6 concludes by discussing some of the implications and suggesting directions for future research.
2 An Equivalence
In this section, I establish the equivalence between OLS predictions and a simplified version of the attention mechanism used in large language models. I start with the standard OLS prediction, then introduce a less conventional but illuminating interpretation–OLS as a similarity-based retrieval process naturally structured within an encoder-decoder framework. I then review the canonical presentation of attention mechanisms as information retrieval, draw a formal connection between the two, and discuss the implications.
2.1 A Proximity Interpretation of Least Squares
Let $\boldsymbol{X}_{\text{train}}\in\mathbb{R}^{N\times P}$ denote the training design matrix and $\boldsymbol{y}\in\mathbb{R}^{N}$ the corresponding response vector, where in-sample observations run from $i=1$ to $N$. The OLS estimator is given by
$\hat{\boldsymbol{\beta}}$ $=(\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}})^{-1}\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{y}.$ (1)
For out-of-sample observations indexed by $j=N+1,\dots,N+J$, the corresponding feature matrix is $\boldsymbol{X}_{\text{test}}\in\mathbb{R}^{J\times P}$, and the predicted values are
$\hat{\boldsymbol{y}}_{\text{test}}$ $=\boldsymbol{X}_{\text{test}}\hat{\boldsymbol{\beta}}$ (2)
$=\boldsymbol{X}_{\text{test}}(\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}})^{-1}\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{y}.$ (3)
What is less widely recognized is that linear regression predictions can be interpreted as a proximity-based estimator. Using the eigendecomposition of $(\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}})^{-1}$, we rewrite the standard out-of-sample prediction formula (2) as
$\hat{\boldsymbol{y}}_{\text{test}}=\boldsymbol{X}_{\text{test}}\boldsymbol{U}\Lambda^{-1}\boldsymbol{U}^{\prime}\boldsymbol{X}_{\text{train}}\boldsymbol{y}_{\text{train}},$ (4)
where $\boldsymbol{U}\Lambda^{-1}\boldsymbol{U}^{\prime}$ is the eigendecomposition of $(\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}})^{-1}$, with $\Lambda$ being the diagonal matrix of eigenvalues of $\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}}$, and $\boldsymbol{U}$ its eigenvectors. Decomposing the $\Lambda$ into $\Lambda^{-\frac{1}{2}}\Lambda^{-\frac{1}{2}}$, we obtain:
$\hat{\boldsymbol{y}}_{\text{test}}=\underbrace{\boldsymbol{X}_{\text{test}}\boldsymbol{U}\Lambda^{-\frac{1}{2}}}_{\boldsymbol{F}_{\text{test}}}\underbrace{\Lambda^{-\frac{1}{2}}\boldsymbol{U}^{\prime}\boldsymbol{X}_{\text{train}}^{\prime}}_{\boldsymbol{F}_{\text{train}}^{\prime}}\boldsymbol{y}_{\text{train}}.$ (5)
This formulation highlights a natural similarity-based interpretation: both $\boldsymbol{F}_{\text{test}}=\boldsymbol{X}_{\text{test}}\boldsymbol{U}\Lambda^{-\frac{1}{2}}$ and $\boldsymbol{F}_{\text{train}}=\boldsymbol{X}_{\text{train}}\boldsymbol{U}\Lambda^{-\frac{1}{2}}$ are standardized factor scores providing an orthonormal representation of $\boldsymbol{X}_{\text{test}}$ and $\boldsymbol{X}_{\text{train}}$, respectively. The representation preserves the original input dimension, so $\boldsymbol{F}_{\text{train}}\in\mathbb{R}^{N\times P}$ and $\boldsymbol{F}_{\text{test}}\in\mathbb{R}^{J\times P}$. For $\boldsymbol{F}_{\text{train}}$, we have that
$\boldsymbol{F}_{\text{train}}^{\prime}\boldsymbol{F}_{\text{train}}=I_{P}.$ (6)
However, since $\Lambda$ and $\boldsymbol{U}$ are estimated from the training data only, the identity $\boldsymbol{F}_{\text{test}}^{\prime}\boldsymbol{F}_{\text{test}}={I}_{P}$ does not hold exactly—just as $\hat{\boldsymbol{\beta}}$ minimizes squared residuals by construction on the training data, but not necessarily on the test sample.
Thus, OLS out-of-sample predictions can be rewritten as:
$\underbrace{\hat{\boldsymbol{y}}_{\text{test}}}_{1\times J}=\underbrace{\left[\boldsymbol{F}_{\text{test}}\boldsymbol{F}_{\text{train}}^{\prime}\right]}_{J\times N}\,\,\underbrace{\boldsymbol{y}_{\text{train}}}_{N\times 1},$ (7)
where $\boldsymbol{F}_{\text{test}}\boldsymbol{F}_{\text{train}}^{\prime}$ is a matrix of inner products between test observations (represented in $\mathbb{R}^{P}$ by $\boldsymbol{F}_{\text{test}}$) and training observations (represented by $\boldsymbol{F}_{\text{train}}$). This reveals OLS as an inner-product-based similarity method, reinforcing its connection to traditional similarity-based estimation such as nearest neighbors. That is, the term $\boldsymbol{F}_{\text{test}}\boldsymbol{F}_{\text{train}}^{\prime}$ serves as a proximity matrix, measuring the similarity between pairs of training and test observations. Unlike the usual dual formula for ridge regression or likewise regularized problems, this view integrates the equivalent of dual coefficients $\hat{\boldsymbol{\alpha}}=(\boldsymbol{X}_{\text{train}}\boldsymbol{X}_{\text{train}}^{\prime}+\lambda I_{N})^{-1}$ directly into the interpretation of test set predictions.
Inner Product and Cosine Similarity Weighting.
The view of OLS as a similarity-based method becomes more evident when considering the equivalent formulation of equation (7) for a single prediction of test sample observation $j$:
$\hat{y}_{j}$ $=\sum_{i=1}^{N}\underbrace{\langle F_{j},F_{i}\rangle}_{\equiv\omega_{ji}}y_{i}.$ (8)
Here, $\langle\cdot,\cdot\rangle$ denotes the Euclidean inner product. OLS can thus be interpreted as a weighted averaging procedure, where the predicted outcome $\hat{y}_{j}$ is a sum of observed responses $y_{i}$, weighted by $\omega_{ji}$, which reflects the similarity between the corresponding factor $F_{i}$ and $F_{j}$ in the Euclidean sense. dual discuss this property as a way to interpret the weights $\omega_{ji}$ in high-dimensional and nonlinear machine learning-based macroeconomic forecasts.
Note that $\langle\cdot,\cdot\rangle$ is an unscaled measure of vector alignment. Indeed, $\omega_{ji}$ can be factorized as
$\hat{y}_{j}$ $=\sum_{i=1}^{N}\underbrace{\|F_{j}\|\|F_{i}\|}_{\text{scale}}\underbrace{\cos(\theta_{ji})}_{\text{alignment}}y_{i},$ (9)
where $\cos(\theta_{ji})=\frac{\langle F_{j},F_{i}\rangle}{\|F_{j}\|\|F_{i}\|}$, $\theta_{ji}$ is the angle between $F_{j}$ and $F_{i}$, and $\|\cdot\|$ denotes the vector norm. We see that large values of $\omega_{ji}$ can arise for two reasons: either because the vectors are highly aligned (large $\cos(\theta_{ji})$) or because observation $i$ has a large norm, meaning it is far from the origin in $\mathbb{R}^{P}$. The largest $\omega_{ji}$ values feature both. It is also worth noting that the term $\|F_{j}\|$ does not affect the relative weighting of observations for a given target $j$, as it is common to all $\omega_{ji}$’s.
Representations similar to that of equation (7) have previously appeared in textbooks, typically in the context of fitted values (i.e., when $\boldsymbol{X}_{\text{test}}=\boldsymbol{X}_{\text{train}}$), as they naturally follow from fundamental matrix properties. Notably, hastie2009elements present the singular-value decomposition of the “hat matrix” as a computational tool or to illustrate the effects of ridge-like shrinkage on the least squares projection matrix. The novelty here lies in reinterpreting this result to connect OLS with proximity-based methods, where proximity is defined via simple inner products in a transformed space. This perspective will later prove useful in establishing links to the attention mechanism, which is grounded in the notion of similarity.
Moreover, it is worth stressing that the factor structure arising from the OLS fit is entirely mechanical. That is, it is a purely algebraic reorganization that holds for any dataset, regardless of its statistical properties. No assumptions about distribution, stationarity, or noise are required, nor does this imply dimensionality reduction. These factors are intrinsic to the OLS solution and exist independently of whether an underlying “true” factor structure is present. However, assigning meaningful interpretations to these factors—such as plotting a column of $\boldsymbol{F}_{\text{train}}$ and linking it to latent scientific concepts—requires additional statistical assumptions, as discussed in, for instance, bai2008large.
On the Choice of the Spectral Decomposition.
While there are multiple ways to decompose a symmetric positive definite matrix, the spectral decomposition is particularly useful for understanding OLS as a similarity-matching system in an uncorrelated $P$-dimensional space. When characteristics are correlated, aggregating similarity scores across dimensions is not straightforward. Using
$\hat{y}_{j}$ $=\sum_{i=1}^{N}\langle X_{j},X_{i}\rangle y_{i}$ (10)
instead of equation (8), even when $\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}}\neq I_{P}$, is suboptimal because it fails to account for covariances between predictors in computing similarities. For example, if two predictors are highly correlated with $|\rho|\approx 1$, equation (10) effectively counts the redundant information twice, disproportionally affecting the estimated prediction weights.
However, in an orthonormal basis where characteristics are mutually uncorrelated, the total similarity between observations $i$ and $j$ can be computed simply as the sum of similarities for each characteristics calculated separately, as in equation (8). This is analogous to standard results on variance calculations: when observations are mutually-uncorrelated, the variance of a sum equals the sum of individual variances. By the same logic, the total similarity in this transformed orthonormal space is just the sum of individual characteristics similarities.
2.2 An Encoder-Decoder View of the Proximity Formulation
More generally, we can define two mappings,
$\boldsymbol{W}_{\text{train}}:\mathbb{R}^{P}\to\mathbb{R}^{P}\quad\quad\text{and}\quad\quad\boldsymbol{W}_{\text{test}}:\mathbb{R}^{P}\to\mathbb{R}^{P},$ (11)
which embed $\boldsymbol{X}_{\text{train}}$ and $\boldsymbol{X}_{\text{test}}$ into a space where their similarity can be directly measured via Euclidean inner products. Specifically, in the case of the OLS solution, these mappings take the form
${\boldsymbol{W}}_{\text{train}}={\boldsymbol{W}}_{\text{test}}=\boldsymbol{U}\Lambda^{-\frac{1}{2}},$ (12)
so that
$\boldsymbol{F}_{\text{train}}=\boldsymbol{X}_{\text{train}}{\boldsymbol{W}}\quad\quad\text{and}\quad\quad\boldsymbol{F}_{\text{test}}=\boldsymbol{X}_{\text{test}}{\boldsymbol{W}}.$ (13)
These transformations define what are often referred to as encoding and decoding operations. The encoding step transforms the original predictors into the factor space, where relationships among predictors are orthogonalized for easier comparison. The decoding step ensures that the same transformation applied to the training set is also applied to the test set, preserving structural alignment and relying solely on mappings parameterized by in-sample data.
In summary, we have shown that, beyond its standard interpretation, the least squares solution also reformulates the prediction problem as a similarity-matching framework, where outcomes $\boldsymbol{y}_{\text{train}}$ are weighted according to inner products of encoded representations of predictors. As is already apparent, the OLS encoding is fairly passive from a modeling standpoint, since it is neither compressing nor expanding the information available in $\boldsymbol{X}$. Indeed, $\boldsymbol{W}_{\text{train}}$ and $\boldsymbol{W}_{\text{test}}$ still map into a space of the same dimension as the original inputs. Still, it is interesting to note that, despite lacking a factorization pre-processing step or an explicit encoding layer, OLS inherently performs orthogonal encoding and decoding operations in the background. In Section 3.1, we will examine $\boldsymbol{W}$’s that induce dimensionality reduction, further aligning the framework with traditional factor models and even autoencoders.
2.3 A Review of the Attention Mechanism
The attention mechanism is a fundamental component of modern neural architectures, particularly in sequence modeling and large language models. First introduced in the context of neural machine translation (bahdanau2014neural), attention mechanisms have since evolved into more sophisticated forms, notably self-attention as popularized by the Transformer architecture (vaswani2017attention). For detailed treatments, in a language more amicable to researchers in economics and finance, I refer the reader to these slides and guijarro2021deep.
At its core, attention dynamically reweights input elements based on similarity scores, allowing models to selectively focus on the most relevant components of an input sequence. The standard attention mechanism is defined as
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\prime}}{\sqrt{P}}\right)\mathbf{V},$ (14)
where:
  • $\mathbf{Q}$ (queries) represents a set of input vectors that seek relevant information.
  • $\mathbf{K}$ (keys) represents the stored information to be retrieved.
  • $\mathbf{V}$ (values) contains the actual content to be weighted and aggregated.
The similarity between queries and keys is computed as a scaled dot product, normalized by the square root of the feature dimension $P$ to stabilize gradients. The softmax function, defined as
$\text{softmax}(z_{i})=\frac{e^{z_{i}}}{\sum_{\tau=1}^{N}e^{z_{\tau}}},$ (15)
where $z_{i}$ represents inputs, ensures that the resulting attention weights form a weighted sum of the values $\mathbf{V}$. This (logistic) transformation effectively acts as a probability distribution over the input elements.
Dot product attention can be viewed as a similarity-based retrieval mechanism that selects and aggregates relevant information from the value matrix based on query-key alignments. The underlying principle is analogous to classical nearest neighbor methods, where similarity measures dictate the weighting of known observations in making predictions. Unlike hard selection techniques such as k-nearest neighbors, attention mechanisms employ soft weighting through the softmax function, allowing gradients to propagate smoothly and enabling end-to-end differentiability needed for training through backpropagation (deeplearning). The scaling factor $\sqrt{P}$ stabilizes the distribution of attention weights by preventing the dot product scores from growing excessively with increasing $P$. Without this adjustment, the softmax outputs could become overly peaked, potentially causing gradients to vanish or explode.
The dominant interpretation of attention, obvious from its mathematical formulation, is that of an information retrieval mechanism. The query $\mathbf{Q}$ represents a search term, the keys $\mathbf{K}$ define a database of stored representations, and the values $\mathbf{V}$ contain the actual content retrieved based on the relevance scores.
Attention as a Kernel Method.
Recent work has shown that attention mechanisms can be understood through the lens of kernel methods, where attention weights correspond to similarity measures in a latent space (tsai2019transformer; choromanski2020rethinking; katharopoulos2020transformers; peng2021random). The attention function can be expressed in terms of kernelized similarity computations:
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\boldsymbol{\gamma}\mathbf{V},\quad\text{where}\quad\gamma_{i\tau}=\frac{k(\mathbf{q}_{i},\mathbf{k}_{\tau})}{\sum_{\tau=1}^{N}k(\mathbf{q}_{i},\mathbf{k}_{\tau})}$ (16)
for some kernel function $k(\cdot,\cdot)$, often instantiated as the exponential of a scaled dot product. This formulation highlights the connection between attention-based learning and classical similarity-based methods such as nearest neighbors and kernel-based methods in general. However, these discussions often do not explicitly connect to a regression-based perspective. However, recalling the connection between infinite-dimensional basis expansions (primal solution) and kernel ridge regression in its dual formulation, it seems natural to expect that a link exists under some conditions. Along those lines, garnelo2023exploring have drawn a more direct parallel between attention mechanisms and regression models, highlighting that regularized least squares solutions can be retreived by a certain generalization of Attention.
Other key topics relevant for the Transformer architecture are multi-head attentions and self-attention. Those and their implications will be discussed after establishing OLS as simplified attention module.
2.4 The OLS-Attention Nexus
OLS predictions for an out-of-sample observation can be expressed through an attention-like formulation using the standard decomposition of linear mappings. Using the encoder-decoder formulation of Section 2.2, the OLS prediction can be rewritten in an attention-like form by introducing matrices $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ analogous to those in the attention mechanism:
$\mathbf{Q}$ $=\boldsymbol{X}_{\text{test}}\boldsymbol{W}_{Q},$
$\mathbf{K}$ $=\boldsymbol{X}_{\text{train}}\boldsymbol{W}_{K},$
$\mathbf{V}$ $=\boldsymbol{y}\cdot\boldsymbol{w}_{v}.$
To generalize this formulation, I replace the softmax function with a general function $g(\cdot)$ and substitute the expressions for $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$:
$\text{Attention}(\boldsymbol{X}_{\text{test}},\boldsymbol{X}_{\text{train}},\mathbf{y};\mathbf{W})=g\left(\boldsymbol{X}_{\text{test}}\boldsymbol{W}_{Q}\boldsymbol{W}_{K}^{\prime}\boldsymbol{X}_{\text{train}}^{\prime}\right)\boldsymbol{y}\cdot\boldsymbol{w}_{v}.$
Now, setting $g(\cdot)=I(\cdot)$ (the identity function), $\boldsymbol{W}_{Q}\boldsymbol{W}_{K}^{\prime}=(\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}})^{-1}$ or equivalently $\boldsymbol{W}_{K}=\boldsymbol{W}_{Q}=\boldsymbol{U}\Lambda^{-\frac{1}{2}}$ as in equation (20), and $\boldsymbol{w}_{v}=\mathbf{1}$, we obtain:
$\text{Attention}(\boldsymbol{X}_{\text{test}},\boldsymbol{X}_{\text{train}},\mathbf{y};\mathbf{W})$ $=g\left(\boldsymbol{X}_{\text{test}}\boldsymbol{W}_{Q}\boldsymbol{W}_{K}^{\prime}\boldsymbol{X}_{\text{train}}^{\prime}\right)\boldsymbol{y}\cdot\boldsymbol{w}_{v}$ (17)
$=\boldsymbol{X}_{\text{test}}(\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}})^{-1}\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{y}$ (18)
$=\boldsymbol{X}_{\text{test}}\hat{\boldsymbol{\beta}}$. (19)
Thus, the out-of-sample OLS prediction can be interpreted as a form of scaled attention, where each test sample attends to the training observations through a similarity measure encoded in an orthonormal $P$-dimensional space.
Using the language of information retrieval systems, in the case of OLS, the query $\mathbf{Q}$ corresponds to $\boldsymbol{X}_{\text{test}}$, encoded with $\boldsymbol{W}_{Q}$ to enable efficient identification of relevant rows of the existing database. To do so, these are matched against keys $\mathbf{K}$, which encode $\boldsymbol{X}_{\text{train}}$ through the same mapping. The entries in the database are the observations $1,\dots,N$, each associated with realized values $\mathbf{V}$ of the dependent variable, which we must linearly combine in order to generate a prediction.
The term cross-attention is often used to describe a sequence attending to another. This characterizes the case where we set $\mathbf{Q}=\boldsymbol{X}_{\text{test}}\boldsymbol{W}_{Q}$, as $\boldsymbol{K}$ is constructed from a different input. By contrast, replacing $\mathbf{Q}=\boldsymbol{X}_{\text{train}}\boldsymbol{W}_{Q}$ while maintaining the same solution for $\boldsymbol{W}_{Q}\boldsymbol{W}_{K}^{\prime}$ yields traditional in-sample fitted values. By sharing a common input for both $\boldsymbol{K}$ and $\mathbf{Q}$ the in-sample predictors setup is closer to self-attention. However, self-attention typically implies a time-series structure, which I address separately in Section 4.2.
As shown in Section 3.2, reintroducing a nonlinear function $g(\cdot)$ maintains a regression interpretation. This is also an opportunity to examine the additional liberties and constraints that arise when transitioning from OLS to a nonlinear Attention Regression.
The OLS Solution as an Optimal Embedding.
A natural question is whether using the plug-in solution $\boldsymbol{W}_{Q}\boldsymbol{W}_{K}^{\prime}=(\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}})^{-1}$ is consistent with first order conditions from a well-defined linear attention problem with $\mathbf{V}=\boldsymbol{y}$. The answer to this is yes.
The standard interpretation of least squares regression involves solving for the optimal coefficient vector, but an alternative formulation seeks to optimize the embedding transformation itself. Specifically, rather than minimizing over $\boldsymbol{\beta}$, we consider the objective
$\min_{\boldsymbol{\Omega}\in\mathbb{R}^{{P}\times P}}\|\boldsymbol{y}-\boldsymbol{X}_{\text{train}}\boldsymbol{\Omega}\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{y}\|^{2},$ (20)
where $\boldsymbol{\Omega}\equiv\boldsymbol{W}_{Q}\boldsymbol{W}_{K}^{\prime}$ represents the transformation of the input space being optimized. The problem in equation (20) is to find the optimal latent space in which to compare observations via inner products, with the objective of creating a weighting scheme that minimizes squared prediction errors. Taking first-order conditions with respect to $\boldsymbol{\Omega}$ and solving the resulting system, we (inevitably) obtain the closed-form solution
$\hat{\boldsymbol{\Omega}}=(\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}})^{-1}.$ (21)
Thus, rather than directly solving for regression coefficients, this formulation reveals that the optimal encoding-decoding scheme to calculate inner product proximity scores is given by the inverse of the sample covariance matrix.
While conceptually appealing, estimating $\boldsymbol{\Omega}$ directly as in problem (20) is not computationally advantageous. The parameter space expands substantially—even when using a factorized form like $\boldsymbol{W}\boldsymbol{W}$’, where $\boldsymbol{W}$ is constrained to be lower triangular, à la Cholesky, to ensure the positive definiteness of $\boldsymbol{W}\boldsymbol{W}$’. Instead of estimating $P$ regression coefficients, we must now estimate $\frac{P(P+1)}{2}$ free parameters. Thus, while the approach offers conceptual insight, especially in terms of understanding regression through the lens of learned inner product spaces, its practical utility in the linear case remains limited. That said, the situation may be different in nonlinear settings.
Putting it all together.
Figure 1 summarizes the correspondence between OLS predictions and the attention mechanism. The test observations $\boldsymbol{X}_{\text{test}}$ and training observations $\boldsymbol{X}_{\text{train}}$ are projected into an orthonormal space via the shared encoding $\boldsymbol{W}=\boldsymbol{U}\Lambda^{-1/2}$, yielding queries $\mathbf{Q}$ and keys $\mathbf{K}$, respectively. The inner product $\mathbf{Q}\mathbf{K}^{\prime}$ produces a similarity matrix whose entries $\omega_{ji}$ weight the training outcomes $\boldsymbol{y}$ to form predictions. In the linear case, these weights are unbounded and may take negative values—properties that distinguish OLS from softmax-based attention, where weights are constrained to the simplex. The matrix $\boldsymbol{\Omega}=\boldsymbol{W}_{Q}\boldsymbol{W}_{K}^{\prime}=(\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}})^{-1}$ emerges as the solution to an optimal embedding problem: it defines the latent geometry in which proximity-based weighting minimizes squared prediction error.
Figure 1: OLS Predictions as a Linear Attention Module
3 Beyond OLS
The equivalence between the attention module and OLS relies on several simplifications with respect to what is actually going on under the hood of Transformers. Adding those back and starting from OLS provides valuable insights. I begin by discussing modifications to the attention module itself, followed by its integration within the broader Transformer architecture.
3.1 Dimensionality Reduction
The OLS solution reveals that the encoding and decoding matrices, $\boldsymbol{W}_{K}$ and $\boldsymbol{W}_{Q}$, serve solely to project $\boldsymbol{X}$ into a space where features are orthogonal, facilitating direct computation of Euclidean inner products. However, in Transformer-based models, the learned mappings $\boldsymbol{W}_{K}$ and $\boldsymbol{W}_{Q}$ typically project data into a lower-dimensional space. This is particularly relevant, as tokens (i.e., words) are usually mapped into a larger conceptual space at the entrance/exit of Transformers networks, before these outputs arrive at the attention module.
How does the OLS-based attention mechanism compare to the learned embeddings in Transformers? The answer lies in Principal Component Regression (PCR). Suppose we redefine
$\boldsymbol{W}_{\text{train}}:\mathbb{R}^{P}\to\mathbb{R}^{L}\quad\quad\text{and}\quad\quad\boldsymbol{W}_{\text{test}}:\mathbb{R}^{P}\to\mathbb{R}^{L},$ (22)
as mappings from a space of dimension $P$ to a lower-dimensional space $L$, where $L<P$, effectively compressing the data representation. In this case, the resulting formulation is closely related to the application of PCA to reduce the dimensionality of the predictor space before applying OLS.
Specifically, consider truncating the eigenvalue decomposition of the sample covariance matrix:
$\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}}=\boldsymbol{U}\Lambda\boldsymbol{U}^{\prime},$ (23)
where, instead of inverting $\Lambda$ directly, we define a low-rank approximation by retaining only the first $L$ principal components, setting the smallest eigenvalues to zero. This corresponds to constructing $\boldsymbol{W}_{K}$ and $\boldsymbol{W}_{Q}$ such that they select the leading eigenvectors:
$\boldsymbol{W}_{K}=\boldsymbol{W}_{Q}=\boldsymbol{U}_{L}\Lambda_{L}^{-\frac{1}{2}},$
where $\Lambda_{L}\in\mathbb{R}^{L\times L}$ is a diagonal matrix containing the first $L$ eigenvalues, and $\boldsymbol{U}_{L}\in\mathbb{R}^{P\times L}$ holds the corresponding $L$ eigenvectors. Therefore, we have $\boldsymbol{F}_{\text{test}}^{L}=\boldsymbol{X}_{\text{test}}\boldsymbol{W}_{Q}$ and $\boldsymbol{F}_{\text{train}}^{L}=\boldsymbol{X}_{\text{train}}\boldsymbol{W}_{K}$, both being lower-dimensional representations of the original data projected onto the subspace spanned by the first $L$ principal components.
In this formulation, it is straightforward to see that PCR’s predictions–essentially performing OLS on $\boldsymbol{F}_{\text{train}}^{L}$, the top $L$ principal components of the input data–reduce to our linear attention mechanism with dimensionality reduction:
$\hat{\boldsymbol{y}}_{\text{test}}$ $=\boldsymbol{F}_{\text{test}}^{L}\hat{\boldsymbol{\theta}}=\boldsymbol{F}_{\text{test}}^{L}\underbrace{\left(\boldsymbol{F}_{\text{train}}^{L^{\prime}}\boldsymbol{F}_{\text{train}}^{L}\right)^{-1}}_{=\boldsymbol{I}_{L}}\boldsymbol{F}_{\text{train}}^{L^{\prime}}\boldsymbol{y}_{\text{train}}$ (24)
$=\boldsymbol{F}_{\text{test}}^{L}\boldsymbol{F}_{\text{train}}^{L^{\prime}}\boldsymbol{y}_{\text{train}}.$ (25)
It is noteworthy that this solution also naturally arises in the linear embedding optimization problem
$\min_{\boldsymbol{\Omega}\in\mathbb{R}^{{P}\times P}}\|\boldsymbol{y}-\boldsymbol{X}_{\text{train}}\boldsymbol{\Omega}\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{y}\|^{2}\quad\quad\text{such that}\quad\quad\text{rank}(\boldsymbol{\Omega})=L\,$ (26)
with $\hat{\boldsymbol{\Omega}}=(\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}})^{-1}$ if $L=P$. To satisfy the rank constraint $\text{rank}(\boldsymbol{\Omega})=L<P$, we must approximate $\boldsymbol{\Omega}$ by a rank-$L$ matrix. From matrix approximation theory (Eckart-Young theorem), the best rank-$L$ approximation in the Frobenius norm is obtained by truncating the smallest eigenvalues of $\boldsymbol{\Omega}$. This corresponds to selecting only the top $L$ eigenvectors of $\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}}$, leading to:
$\hat{\boldsymbol{\Omega}}_{L}=\boldsymbol{U}_{L}\Lambda_{L}^{-1}\boldsymbol{U}_{L}^{\prime}.$ (27)
This follows from low-rank approximation theory, where truncating small eigenvalues minimizes the reconstruction error. The matrix $\hat{\boldsymbol{\Omega}}_{L}$ retains only the most significant principal components, effectively performing a dimension reduction while minimizing information loss.
3.2 Nonlinearity and the Use of Softmax
Reintroducing nonlinearity—along with the dimension reduction discussed above—leads us to the full attention mechanism used in Transformer architectures. The row-wise application of the softmax function $g(\cdot)$ brings nonlinearities into the attention-based regression framework, making it impossible to derive $\boldsymbol{\Omega}$ in closed form from
$\min_{\boldsymbol{\Omega}\in\mathbb{R}^{P\times P}}\|\boldsymbol{y}-\text{softmax}\left(\boldsymbol{X}_{\text{train}}\boldsymbol{\Omega}\boldsymbol{X}_{\text{train}}^{\prime}\right)\boldsymbol{y}\|^{2}\,.$ (28)
A key distinction of $\text{softmax}()$ outputs—akin to probabilities derived from a logistic likelihood function—is that the resulting weights on outcomes are strictly positive and sum to one. Thus, this nonlinear transformation also implies restrictions on $y_{i}$’s may be used to predict $y_{j}$.
Notably, this means that predictions are constructed solely from non-negative weights on $\boldsymbol{y}$ that sum to one, preventing the model from assigning negative weights even when $\boldsymbol{X}_{\text{train}}$ and $\boldsymbol{X}_{\text{test}}$ lie on opposite ends in $\mathbb{R}^{P}$. Large negative entries in $\boldsymbol{X}_{\text{train}}\boldsymbol{\Omega}\boldsymbol{X}_{\text{train}}^{\prime}$, when passed through the exponential function in $\text{softmax}()$, effectively yield weights of approximately zero. The same behavior is observed in Random Forest, where predictions are constructed using non-negative weights on observed responses, as discussed in dual.
In contrast, linear regression with $g(\cdot)=I(\cdot)$ allows both positive and negative weights on the target variable. While the non-negative restriction makes sense in language models — where attention acts on transformed tokenized data — it may prove limiting in other applications where the ability to assign negative weights and exploit symmetry is useful. When $\text{softmax}()$ is applied without multi-head attention and attention modules are recursively chained, this constraint sticks, introducing nonlinearity while simultaneously preventing the model from leveraging symmetrical information in the data. That is, there is no such thing as negative attention. While this may seem reasonable from a behavioral standpoint, OLS—along with a slate of more sophisticated nonlinear methods—makes use of negative attention all the time.
Any Use for an Attention Regression?
Given the many uses of OLS in statistical practice, an empiricist may rightfully wonder at this point: is there any empirical value in running nonlinear Attention Regressions such as the one outlined in equation (28)? The answer is: perhaps.
Equation (28) represents a nonlinear regression model with a simplex constraint on proximity weights (that is, the analog to $\omega_{ji}$ in equation (8)). The role of the $\text{softmax}()$ function is not merely to zero out the negative values of a predefined solution; rather, the optimized $\hat{\boldsymbol{\Omega}}$ will generally differ from the OLS solution, $(\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}})^{-1}$. Nevertheless, regardless of the form $\hat{\boldsymbol{\Omega}}$ ultimately takes, the resulting solution still implies that observations are compared in a space that remains linearly mapped to the original one, and then resulting values are passed through a nonlinear transformation.
This observation brings us back to equation (16) and the kernel interpretation of attention. The key distinction between an exponential inner product kernel and the $\text{softmax}()$ function lies in row-wise normalization. While this normalization may initially seem like a defining feature of $\text{softmax}()$ compared to our simpler OLS-based attention mechanism, it is not. In fact, row-wise normalization to 1 is inherent even in OLS. Provided that $\boldsymbol{X}_{\text{train}}$ includes an intercept, all the rows of the projection matrix $\boldsymbol{P}_{\boldsymbol{X}_{\text{train}}}$ sum to 1. Consequently, the requirement for attention weights to sum to one offers little distinction from OLS properties: the OLS fitted value for $y_{i}$ is always a linear combination of $\boldsymbol{y}$ with weights summing to 1 — and this holds true for all ii. [^1] By principles of optimality, it is reasonable to expect that solutions to equation (28), when replacing $\text{softmax}()$ with $\text{exp}()$, would still produce rows that approximately sum to 1. This outcome would arise naturally through a different estimate of $\hat{\boldsymbol{\Omega}}$ than the one obtained using $\text{softmax}()$.
The true nonlinear distinction between equation (28) and equation (17) lies in the exponential operation itself. This operation introduces nonlinearity by both squashing negative values toward zero and amplifying large positive ones, while also precluding the use of symmetrical information. Consequently, the Attention Regression framework described in equation (28) can be interpreted as a form of nonparametric regression that (i) retains the row-wise sum-to-one property inherent to the OLS projection matrix and (ii) employs a row-stochastic matrix to map $\boldsymbol{y}$ into $\hat{\boldsymbol{y}}$. However, for that matter, there is nothing sacred about $\text{softmax}()$ and $\text{exp}()$; other $g(\cdot)$’s can also satisfy the same properties. For instance, a normalized ReLU transformation—defined as $\text{ReLU}(x)=\max(0,x)$—could likewise meet conditions (i) and (ii). Alternatively, if one wishes to relax condition (ii) and permit mild negative weighting without resorting to full linearity, a normalized ELU transformation—defined as $\text{ELU}(x)=x$ for $x>0$ and $\nu(e^{x}-1)$ for $x\leq 0$—could be employed instead. While nearly everything is possible in theory, it is worthwhile to keep in mind that smoother transformations lend themselves more easily to gradient-based optimization.
Likely more intriguing for empirical work would be to include equation (28) as the final layer of a standard feedforward neural network. In this setting, predictors would first undergo nonlinear transformations before being used to assess similarities. The simplex constraint on attention weights may prove beneficial in stabilizing the network’s predictions, forcing it to leverage similarities and neglect dissimilarities. This mirrors the advantage seen in Random Forest, which also enforce a form of simplex constraint and are recognized for their capacity to avoid catastrophic prediction errors (MRFjae; goulet2024bag).
In light of this, the Attention Regression emerges as a possibly useful hybrid, combining smooth nonlinearities found in kernel methods and neural networks with the simplex constraint on weights, a feature more commonly seen in tree-based models. Whether this translates into meaningful gains in predictive accuracy is an application-dependent empirical question. Section 5 takes up this question directly, evaluating the finite-sample performance of Attention Regression against standard nonlinear benchmarks across a range of data-generating processes.
Parallels to Multinomial Logistic Regression?
Recalling that the softmax function is effectively the multinomial logit probability function, one might be tempted to draw a connection between equation (28) and multinomial logistic regression. The latter is a regression model that employs a logistic link function to ensure that predicted probabilities fall between 0 and 1 and collectively sum to 1. It is nonlinear in index; that is, nonlinearity arises only through the link function, while the predictors enter the model linearly.
However, in equation (28) above, the softmax is not applied to a single linear prediction but rather to the equivalent of the “hat matrix” $\boldsymbol{A}=\text{softmax}\left(\boldsymbol{X}_{\text{train}}\boldsymbol{\Omega}\boldsymbol{X}_{\text{train}}^{\prime}\right)$, as it is often referred to in machine learning textbooks (hastie2009elements). This ensures that the rows of $\boldsymbol{A}$ resides on the unit simplex, guaranteeing that the final predictions are a convex combination of the elements in $\boldsymbol{y}$. Notably, if $y_{i}\in[0,1]\,\forall i$, this simplex constraint on $\boldsymbol{a}_{i}\equiv\boldsymbol{A}_{\{i,:\}}$ also ensures that $\hat{y}_{i}\in[0,1]\,\forall i$. In this respect, this property mirrors the implication of the logistic link function in multiclass classification problems. But, as we have discussed, it carries more implications, notably, the constraint on the weights $\boldsymbol{a}_{i}$ prevents the use of symmetric information, as would typically occur in a linear model. Hence, if our sole objective is to ensure that $\hat{y}_{i}\in[0,1]\,\forall i$, applying $\text{softmax}()$ directly to the predictions (instead of $\boldsymbol{A}$) achieves this without imposing a non-negativity constraint on the attention weights.
3.3 Regularization and Extending the Attention View to Ridge Regression
Neural network training involves stochastic gradient descent with carefully selected learning rates and many other hyperparameters, a process which implicitly regularizes the model. Thus, it is instructive to return to OLS and its attention-based formulation to consider how explicit regularization alters the embedding interpretation of least squares problems.
As is well-known, the ridge regression solution explicitly regularizes the covariance inversion, leading to the Tikhonov-regularized inverse:
$\hat{\boldsymbol{\Omega}}_{\lambda}=(\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}}+\lambda{I}_{P})^{-1}.$ (29)
This formulation reveals that the usual ridge regularization pushes the estimated embedding mapping toward orthonormality ${I}_{P}$, and preventing overfitting by constraining the influence of small eigenvalues on $\hat{\boldsymbol{\Omega}}_{\lambda}$. In terms of spectral decomposition, the regularized inverse takes the form:
$\hat{\boldsymbol{\Omega}}_{\lambda}=\boldsymbol{U}(\Lambda+\lambda{I}_{P})^{-1}\boldsymbol{U}^{\prime},$ (30)
where $\boldsymbol{U}$ is the eigenvector matrix of $\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}}$, and $\Lambda$ is the diagonal matrix of its eigenvalues. The ridge penalty ensures that the eigenvalues in the neighborhood of 0 are not inverted to excessively large values.
All of this is standard textbook material. However, what is less commonly known is that the resulting $\boldsymbol{F}_{\text{train}}^{\lambda}$ deviates from orthonormality—meaning it no longer fully satisfies the identity $\boldsymbol{F}_{\text{train}}^{\lambda^{\prime}}\boldsymbol{F}_{\text{train}}^{\lambda}={I}_{P}$, as was the case in equation (6). While OLS provides an embedding that perfectly aligns with the inverse covariance structure, ridge regression regularizes this embedding by pushing it towards the prior of an orthogonal basis. In terms of attention mechanisms, this implies that proximity scores are computed as if $\boldsymbol{F}_{\text{train}}^{\lambda^{\prime}}$ were orthogonal, even though this condition no longer holds exactly on the training data. However, this “imperfect” rotation could still be beneficial: it may yield a $\boldsymbol{F}_{\text{train}}^{\lambda}$ that is closer (than OLS’ $\boldsymbol{F}_{\text{train}}$) to the infeasible, ex-post optimal representation $\boldsymbol{F}_{\text{test}}^{*}$, where eigenvectors and eigenvalues would have been estimated directly on the test sample.
Thus, viewing ridge regression through the lens of attention mechanisms provides a different perspective on the bias-variance tradeoff: the embedding transformation may become less representative of the true correlation structure in the data. As a result, similarity computations between observations $i$ and $j$ rely on an assumption of mutual orthogonality in $\boldsymbol{F}_{\text{train}}^{\lambda^{\prime}}$, even though this assumption is violated. The extent of this misalignment depends on the chosen value of the tuning parameter $\lambda$.
4 From Attention Modules to the Transformer
Thus far, the discussion has focused on the attention mechanism itself, treating it as a distinct entity. However, its role within Transformer architectures—where attention is famously “all you need”—involves numerous operations both before data reaches the attention modules and after it passes through them. The original Transformer architecture by vaswani2017attention consists of 6 encoder and 6 decoder layers, each combining multi-head self-attention with feedforward networks, wrapped in residual connections and layer normalization. At each layer, the attention output is added to the input and passed through a feedforward network, with this process recursively repeated across layers.
This section outlines key components, including multi-head and self-attention, pre-trained word embeddings, time series features such as masking, pooled (or global) panel estimation, and the recursive chaining of multiple attention modules.
4.1 Multi-Head Attention
An important component of the transformer architecture is multi-head attention, which enables multiple attention mechanisms to operate in parallel, each with distinct $\boldsymbol{W}_{K}$, $\boldsymbol{W}_{Q}$, and $\boldsymbol{W}_{V}$ matrices. It is defined as:
$\text{MultiHead}(\mathbf{Q},\mathbf{K},\mathbf{V})$ $=\text{Concat}(\text{head}_{1},\dots,\text{head}_{H})\mathbf{W}_{O},$ (31)
$\text{where}\quad\text{head}_{h}$ $=\text{Attention}(\mathbf{Q}\mathbf{W}_{h}^{Q},\mathbf{K}\mathbf{W}_{h}^{K},\mathbf{V}\mathbf{W}_{h}^{V}).$ (32)
Each attention head applies a separate linear transformation to the inputs, allowing the model to learn different attention patterns across subspaces of the feature space. This facilitates diverse contextualization of words, leading to different embeddings and alternative forms of dimensionality reduction.
In the linear case without dimensionality reduction in equation (20), it is evident from the optimization problem that these multiple heads would be redundant, ultimately collapsing to the solution for a single head.
In the nonlinear case, the situation changes significantly. The same rationale that justifies wide neural networks applies here as well, where multiple nonlinear transformations of the inputs can be learned and effectively combined through supervision. This can also been seen as some form of stacking (wolpert1992stacked) or regression-based forecasts combination (granger1984combining). Moreover, the nonlinear case requires numerical optimization rather than closed-form solutions, introducing randomness through parameter initialization. Combining multiple heads helps mitigate this uncertainty.
4.2 Self-Attention, Word2Vec, and Vector Autoregressions
Thus far, the time series or sequential aspects have not been directly addressed. In self-attention, the same set of embeddings serves as queries, keys, and values, allowing each element in the sequence to weigh all others in computing its representation. Self-attention is particularly effective in capturing long-range dependencies in sequences, a key limitation of the previous generation of models that used recurrent architectures. While the latter is more akin to state-space models, the attention framework aligns more naturally with traditional (vector) autoregression estimated through least squares. In time series econometrics, it is well known that extraction of latent states can be performed through traditional filtering tools or via direct projection-based extractions, as emphasized in, e.g., hamilton2018you.
This tension between filtering-based approaches and projections on lagged values also arises in the AI literature. While attention mechanisms can handle long sequences statistically, their computational cost scales quadratically with sequence length due to the $T\times T$ attention matrix. This makes standard self-attention, as used in Transformers, a bottleneck for long sequences and has motivated a return to more classical time series ideas — such as the structured state-space approach of gu2023mamba or the use of fast Fourier transforms in fein2025fft. In this light, interpreting attention as a form of latent-state estimation via regression becomes quite natural.
Univariate Autoregression.
Consider an autoregressive model of order 1 (AR(1)), where we focus solely on modeling in-sample dynamics. OLS fitted values take the form:
$\hat{\boldsymbol{y}}=\boldsymbol{a}\boldsymbol{y},$ (33)
where $\boldsymbol{a}=\boldsymbol{y}_{-1}(\boldsymbol{y}_{-1}^{\prime}\boldsymbol{y}_{-1})^{-1}\boldsymbol{y}_{-1}^{\prime}$, which is approximately numerically equivalent to $\boldsymbol{a}=\boldsymbol{y}_{-1}(\boldsymbol{y}^{\prime}\boldsymbol{y})^{-1}\boldsymbol{y}_{-1}^{\prime}$ under stationarity and with a sufficiently long time series. Intuitively, the temporal covariance structure is the same for $\boldsymbol{y}$ and $\boldsymbol{y}_{-1}$, as they represent the same series with a shifted time index. Since $\boldsymbol{y}_{-1}$ and $\boldsymbol{y}$ nearly perfectly overlap, the sums $\boldsymbol{y}_{-1}^{\prime}\boldsymbol{y}_{-1}$ and $\boldsymbol{y}^{\prime}\boldsymbol{y}$ differ only by a single observation. In this view, the attention weights for $\boldsymbol{y}$, which are $\boldsymbol{a}=(L\boldsymbol{y})(\boldsymbol{y}^{\prime}\boldsymbol{y})^{-1}(L\boldsymbol{y}^{\prime})$ using the lag operator ($L$) notation, are inherently a function of the series itself.
Word2Vec and Pre-trained Embeddings.
If $\boldsymbol{y}$ were a sequence of words, they would enter the network architecture as $\boldsymbol{Y}\in\mathbb{R}^{N\times M}$, where $N$ remains the length of the sequence and $M$ represents the size of its alternative representation, typically obtained through procedures such as Word2Vec. These pre-trained word embeddings assign each word an $M$-dimensional continuous vector that captures semantic and syntactic relationships based on word co-occurrence patterns in large corpus.
Word2Vec, a shallow neural network trained to predict context (in the skip-gram model) or target words (in the CBOW model) (mikolov2013efficient), has been shown to behave similarly to matrix factorization, closely approximating low-rank or SVD-like decompositions of word co-occurrence matrices (levy2015improving). In this light, Word2Vec can be viewed as a compressed or nonlinear regularized factor model applied to language data. In light of the discussion of OLS in Section 2.1, the association between problems framed as predictive tasks and those involving factor extraction is no mere coincidence: even OLS implicitly performs matrix factorization to generate predictions based on similarity.
How can such operations be linked to linear time series modeling trained with least squares?
Vector Autoregression.
Consider $\boldsymbol{y}$ as a sequence of a given variable, such as inflation, where the goal is to determine which past values should be weighted to produce the most accurate forecast of the next realization. Using $\boldsymbol{y}$ in an autoregressive (AR($p$)) model provides only a limited notion of context. However, an economy at time $\tau$ can be characterized in a much richer space than what is captured by past values of inflation. When comparing economic conditions at time $\tau$ with those at time $t$, analysts typically consider a broader set of indicators beyond inflation, including interest rates, GDP growth, and unemployment. In this setting, the matrix of vectorized representations of our “inflation tokens” takes the form $\boldsymbol{Y}\in\mathbb{R}^{N\times M}$, where $M$ corresponds to the number of macroeconomic series used to characterize the context. In the spirit of Word2Vec and related approaches, these $M$ series could be augmented or replaced by factors pre-extracted from a large panel of macroeconomic indicators.
Incorporating this structure into the framework of equation (33), we obtain
$\hat{\boldsymbol{Y}}=\boldsymbol{A}\boldsymbol{Y},$
where
$\boldsymbol{A}=\boldsymbol{Y}_{-1}(\boldsymbol{Y}_{-1}^{\prime}\boldsymbol{Y}_{-1})^{-1}\boldsymbol{Y}_{-1}^{\prime},$
which is numerically equivalent to
$\boldsymbol{A}=(L\boldsymbol{Y})(\boldsymbol{Y}^{\prime}\boldsymbol{Y})^{-1}(L\boldsymbol{Y})^{\prime}$
for the same reasons discussed earlier. This clearly corresponds to the fitted values of a vector autoregression (VAR), where the in-sample predictors of the $M$ variables at each time $t$ can be interpreted as the context vector. In macroeconomic modeling, this context vector can be understood as model-consistent expectations of the variables entering the system (sims1980macroeconomics). In the natural language processing environment, this can be interpreted as sentence-consistent (or more generally, sequence-consistent) notions of context for each word.
4.3 Masking and Pseudo-Out-of-Sample Evaluation
Another interesting component of the application of attention in the decoder part of certain LLMs is the use of masking, which ensures that predictions at time $t$ do not have access to future information from $t+1$ onward. This is done by imposing a lower triangular structure on the attention matrix, setting entries corresponding to future time steps to zero. In essence, masking enforces a (Granger) causal structure, yet it does so on $\boldsymbol{a}$, not on $\boldsymbol{\beta}$. This is an interesting perspective from a classical time series econometrics viewpoint, as the OLS-based AR(1) in-sample prediction for observation $t$ is given by:
$\hat{{y}}_{t}=\sum_{\tau=1}^{N}a_{\tau}y_{\tau},$ (34)
which is a weighted average of all the $y_{\tau}$’s in the sequence, whether $y_{\tau}$ is located before or after $y_{t}$. Indeed, while the AR(1) OLS regression predicts the next value in the sequence ($y_{t+1}$) based on an estimated coefficient and $y_{t}$, the proximity/attention view, articulated through $a_{\tau}$, makes it clear that there is some form of information leakage from the future to the past when looking at fitted values. This “leakage” is not apparent from the AR(1) data-generating process since it comes from the estimation of parameters, not the design of the model.
A relevant consideration is whether masking ideas could be implemented to some benefit in traditional time series models. A masking scheme addressing this in our simplified framework takes the form:
$\hat{{y}}_{t}=\frac{1}{\sum_{\tau=1}^{N}a_{\tau}\mathbf{1}(\tau\leq t)}\sum_{\tau=1}^{N}a_{\tau}\mathbf{1}(\tau\leq t)y_{\tau},$ (35)
where $\mathbf{1}(\tau\leq t)$ is an indicator function ensuring that each prediction at time $t$ only attends to previous or current observations. The denominator normalizes the weights to sum to 1, maintaining consistency with linear regression featuring an intercept. In this way, outcomes from the future are excluded from the weighted average designed to explain the past.
This is an intriguing proposition, as researchers in forecasting often rely on backtesting or pseudo-out-of-sample evaluation, recursively re-estimating models to mitigate overfitting and hindsight bias. However, these methods come with significant computational costs, leaving supercomputing clusters worldwide wondering if masking ideas could offer them a much-needed break. Unfortunately, the answer is no: some leakage still persists. The notion of proximity, even when obtained through $(\boldsymbol{y}_{-1}^{\prime}\boldsymbol{y}_{-1})^{-1}$, implies that the similarity between $y_{\tau}$ and $y_{t}$, despite the former preceding the latter, is still assessed through the lens of the $\hat{\boldsymbol{\Omega}}$ matrix, which is estimated using data from $t=1$ to $N$.
To illustrate, consider the task of determining how to up-weight or down-weight March 1974 U.S. inflation data when forecasting February 1978. Using a training sample spanning from the early 1970s to the 2000s, these observations would appear proximate—and rightfully so—as the $\hat{\boldsymbol{\Omega}}$ matrix would reveal that inflation dynamics in the 1970s were distinct from those of the 1990s yet relatively stable within the decade itself. However, if the training sample were limited to data available only up to January 1978, comprising observations from the late 1960s and early 1970s, this proximity would be less apparent. Without the broader historical context including the stable, low-inflation regime of the 1990s and 2000s, March 1974 and February 1978 will appear, in fact, quite distant.
In sum, since proximity is measured in a space defined by a covariance matrix that reflects the entire sample, simply masking future realizations of $y_{t}$ (as in equation (35)) is not enough to replace the widely used pseudo-out-of-sample recursive experiments.
4.4 Links to Pooled Panel Estimation: Positional Encoding and Training
To estimate the densely-parametrized $\boldsymbol{W}_{Q}$, $\boldsymbol{W}_{K}$, and $\boldsymbol{W}_{V}$ for multiple heads—along with the myriad of parameters from other parts of the architecture—a standard time series estimation approach, such as in Section 4.2, is clearly infeasible, even for long sequences. Instead, training requires a large number of sequences, $S$, structured as $\boldsymbol{Y}_{s}\in\mathbb{R}^{N\times M}$. The estimation process then proceeds in the spirit of a pooled panel regression.
While the panel is inherently unbalanced and timestamps do not necessarily align across sequences $s\in\mathcal{S}$, we can define a maximal sequence length (2048 in GPT-3, much longer in GPT-4) and pad shorter sequences with zeros. This setup resembles a standard pooled panel VAR, where positional encoding plays a role analogous to lags in time series models. The key distinction is that lags encode temporal information explicitly in the time domain, whereas positional encoding–typically implemented via thousands of sinusoidal functions–operates in the frequency domain. For simpler VAR models, it is well known that equivalent estimation can be performed in either domain, though time-domain estimation is generally more practical for traditional econometric applications. Once temporal structure has been incorporated via feature engineering, the estimation process treats this as given and focuses on the (numerous) remaining parameters.
While training a pooled panel VAR on such a heterogeneous dataset may seem excessively restrictive–imposing strong homogeneity assumptions on parameters–LLMs embrace the global model approach. Flexibility is restored by leveraging a highly flexible, overparameterized, yet shared conditional mean. This complexity is achieved through the billions of parameters populating the multiple (nonlinear) attention heads, dense feedforward layers processing attention modules’ outputs, and other modules. In contrast, a local model approach would require training separate, tightly constrained models on individual sequences. Increasingly, the global model approach is gaining traction outside of language processing, as it enables the use of complex machine learning algorithms even for moderate-length time series ($N$) by leveraging a large $S$ (hewamalage2022global). A well-known and successful application of this approach in asset pricing is gu2020empirical, which predicts returns using a longitudinal data set that spans 50 years and incorporates multiple characteristics of firms. In the case of LLMs, $\mathcal{S}$ is a curated subset of the entire internet.
To bring everything together, consider the following analogy. Suppose we aim to train a pooled panel VAR to describe the dynamics of multiple economies. We have $M$ variables for $S$ economies, each observed over $N$ months. Temporal information is incorporated by applying the lag operator to each economy’s data matrix, $\boldsymbol{Y}_{s}\in\mathbb{R}^{N\times M}$, for all $s$ and $p$:
$\boldsymbol{Y}_{s,-p}=L^{p}\boldsymbol{Y}_{s}.$
We then concatenate the transformed data appropriately and estimate the system using OLS after removing fixed effects. A linear VAR approach would be highly restrictive, particularly with a heterogeneous panel of diverse economies. To address this, we move beyond linearity and introduce a more sophisticated conditional mean function, allowing for heterogeneity and locality through complex nonlinearities.
In the case of LLMs, the variables correspond to word tokens embedded in a conceptual space of dimension $M$, while the set of economies ($\mathcal{S}$) is analogous to a corpus of text sequences, each of maximal length $N$. Positional encoding injects sequential proximity information into the original word representations, much like lags do in time series modeling. Once all those elements are in place, the challenge lies in feasibly optimizing the immensely complex nonlinear, nonparametric pooled panel regression.
4.5 Recursive Chaining of Attention Modules and Feed-Forward Networks
The original Transformer architecture, introduced by vaswani2017attention, consists of 6 encoder layers and 6 decoder layers, for a total of 12 Transformer blocks. Each encoder block contains a multi-head self-attention mechanism followed by a feedforward neural network, both enclosed within residual connections and layer normalization. That is, the output of the attention mechanism is combined with the original inputs and passed through a standard feedforward neural network. The output of this network is then fed into the next attention module, and this recursive chaining of blocks continues. For a visual representation, see the original diagram in vaswani2017attention or, more recently, the illustration in kelly2025. Decoder blocks follow a similar structure but also include an additional cross-attention layer that attends to the encoder’s output.
From a deep learning perspective, the recursive design of Transformers is intuitive: stacking multiple layers of similar operations has long been a hallmark of success, as seen in convolutional neural networks for computer vision. However, drawing direct analogies between Transformers and traditional statistical models like OLS is more tenuous, since stacking is largely redundant in the case of linear operations. For example, regressing $\boldsymbol{y}_{\text{train}}$ on both the original inputs $\boldsymbol{X}_{\text{train}}$ and the OLS-fitted values $\hat{\boldsymbol{y}}_{\text{train}}$—obtained from a previous run of the same regression—is redundant, since $\hat{\boldsymbol{y}}_{\text{train}}$ lies in the column space of $\boldsymbol{X}_{\text{train}}$ by construction. This follows from the idempotent property of the OLS projection matrix. As a result, beyond the attention mechanism itself, the plain connection to OLS loses traction.
That said, our earlier discussion of latent state extraction via VARs system-consistent expectations suggests a more promising avenue. From this perspective, the encoder and decoder stacks of a Transformer may be viewed as a nonlinear, multi-level state-space system, with attention mechanisms functioning as adaptive, regression-based filters for extracting latent structure.
Nested levels of filtering can also occur in a much more basic setting. Many successful unobserved components models (or structural time series models) also involve nested latent structures (harvey1990forecasting). For example, the well-known local level model is:
$y_{t}$ $=\mu_{t}+\varepsilon_{t},\quad$ $\varepsilon_{t}$ $\sim\mathcal{N}(0,\sigma^{2}_{\varepsilon})$ (36)
$\mu_{t}$ $=\mu_{t-1}+\eta_{t},\quad$ $\eta_{t}$ $\sim\mathcal{N}(0,\sigma^{2}_{\eta})\,.$ (37)
It is conceptually straightforward to expand this into a more sophisticated nested structure. The local linear trend model adds a time-varying slope component:
$y_{t}$ $=\mu_{t}+\varepsilon_{t},\quad$ $\varepsilon_{t}$ $\sim\mathcal{N}(0,\sigma^{2}_{\varepsilon})$ (38)
$\mu_{t}$ $=\mu_{t-1}+\beta_{t-1}+\eta_{t},\quad$ $\eta_{t}$ $\sim\mathcal{N}(0,\sigma^{2}_{\eta})$ (39)
$\beta_{t}$ $=\beta_{t-1}+\zeta_{t},\quad$ $\zeta_{t}$ $\sim\mathcal{N}(0,\sigma^{2}_{\zeta})\,.$ (40)
And this, in turn, can be extended into a local quadratic trend model, introducing a stochastic acceleration term:
$\mu_{t}$ $=\mu_{t-1}+\beta_{t-1}+\tfrac{1}{2}\gamma_{t-1}+\eta_{t},\quad$ $\eta_{t}$ $\sim\mathcal{N}(0,\sigma^{2}_{\eta})$ (41)
$\beta_{t}$ $=\beta_{t-1}+\gamma_{t-1}+\zeta_{t},\quad$ $\zeta_{t}$ $\sim\mathcal{N}(0,\sigma^{2}_{\zeta})$ (42)
$\gamma_{t}$ $=\gamma_{t-1}+\xi_{t},\quad$ $\xi_{t}$ $\sim\mathcal{N}(0,\sigma^{2}_{\xi})\,.$ (43)
In this progression, each level recursively builds upon the filtration of the previous one, much like the deep stacking of layers in Transformers. A key limitation of models such as the local quadratic trend is that, when estimated on a single time series of moderate length, the variance components often collapse toward zero as the number of latent layers increases. Alternatively, the proliferation of latent states—and hence parameters—can lead to overfitting and degraded out-of-sample performance. However, this complexity constraint goes away in the large-$S$ regime, where a vast number of sequences are modeled jointly, as discussed in Section 4.4. It becomes feasible to recursively project and refine latent states across multiple layers.
5 Experimental Evidence
We conduct a Monte Carlo experiment to evaluate the finite-sample properties of the attention-based regression framework discussed in Section 3.2. While I have established that OLS represents a specific case of linear attention, this experiment assesses the performance of the multi-head nonlinear attention regression across a range of functional forms and signal-to-noise environments.
Attention-Based Regression Specification.
I implement the multi-head extension of the nonlinear attention regression framework introduced in Section 3.2. For a matrix of test points $\boldsymbol{X}^{*}$ and training data $\boldsymbol{X}_{\text{train}}$ with outcomes $\boldsymbol{y}$, the estimator forms predictions by aggregating $M$ distinct attention mechanisms:
$\hat{\boldsymbol{y}}=\sum_{m=1}^{M}\alpha_{m}\left[\text{softmax}\left(\boldsymbol{X}^{*}\boldsymbol{\Omega}_{m}\boldsymbol{X}_{\text{train}}^{\prime}\right)\boldsymbol{y}\right],$ (44)
where $\alpha_{m}$ are scalar, head-specific combination weights. Each head learns a unique similarity metric $\boldsymbol{\Omega}_{m}$, which I parameterize via the Cholesky factorization $\boldsymbol{\Omega}_{m}=\boldsymbol{L}_{m}\boldsymbol{L}_{m}^{\prime}$ (with $\boldsymbol{L}_{m}$ being lower triangular). This ensures that $\boldsymbol{\Omega}_{m}$ remains positive semi-definite, effectively mapping the data into a latent space where similarities are computed as valid inner products.
The model parameters $\{\alpha_{m},\boldsymbol{L}_{m}\}_{m=1}^{M}$ are jointly estimated by minimizing the penalized squared error loss:
$\mathcal{L}=\left\|\boldsymbol{y}-\sum_{m=1}^{M}\alpha_{m}\text{softmax}\left(\boldsymbol{X}_{\text{train}}(\boldsymbol{L}_{m}\boldsymbol{L}_{m}^{\prime})\boldsymbol{X}_{\text{train}}^{\prime}\right)\boldsymbol{y}\right\|^{2}+\lambda\sum_{m=1}^{M}\|\boldsymbol{L}_{m}\|_{F}^{2}.$ (45)
Optimization is performed via L-BFGS with early stopping. To anchor the nonconvex optimization near a sensible starting point, each $\boldsymbol{L}_{m}$ initialized using the Cholesky factor of the ridge-regularized precision matrix $(\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}}+\lambda\boldsymbol{I})^{-1}$, scaled by $\sqrt{N}$, with added Gaussian noise to promote diversity across heads.
Data-Generating Processes.
Let $\boldsymbol{X}_{\text{train}}\in\mathbb{R}^{N\times 5}$ with rows drawn independently from $\text{Uniform}[0,1]^{5}$. For each DGP, outcomes are generated according to
$y_{i}=f(\boldsymbol{X}_{i})+\varepsilon_{i},\qquad\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2}),$
where $\sigma^{2}=\text{Var}(f(\boldsymbol{X}_{i}))/\text{SNR}$ calibrates the noise to a target signal-to-noise ratio. I employ six standard regression functions to test the model’s adaptability:
  1. 1.
    Linear: $f(\boldsymbol{X}_{i})=2(X_{i,1}-0.5)-(X_{i,2}-0.5)+3(X_{i,3}-0.5)+1.5(X_{i,4}-0.5)+0.5(X_{i,5}-0.5).$
  2. 2.
    Friedman 1: $f(\boldsymbol{X}_{i})=10\sin(\pi X_{i,1}X_{i,2})+20(X_{i,3}-0.5)^{2}+10X_{i,4}+5X_{i,5}.$
  3. 3.
    Friedman 2: $f(\boldsymbol{X}_{i})=\sin\bigl(\pi(X_{i,1}+X_{i,2}+X_{i,3})\bigr)+\log(1+X_{i,4}^{2}).$
  4. 4.
    Friedman 3: $f(\boldsymbol{X}_{i})=X_{i,1}X_{i,2}+\log(X_{i,3}+X_{i,4}+2).$
  5. 5.
    Rotated Sine: $f(\boldsymbol{X}_{i})=\sin\Bigl(3\sum_{j=1}^{4}X_{i,j}\Bigr).$
  6. 6.
    Soft Radial: $f(\boldsymbol{X}_{i})=\bigl(1+5\|\boldsymbol{X}_{i}-0.5\boldsymbol{1}\|^{2}\bigr)^{-1}.$
These functions span linear, additive nonlinear, interactive, periodic, and radial structures, providing a diverse testbed that does not privilege any specific estimator.
Design and Competitors.
The training sample size is $N\in\{500,1000,2500,5000\}$ and signal-to-noise ratio $\text{SNR}\in\{0.5,1,2,3\}$. A fixed test set of $J=1000$ observations is used throughout, with results averaged over 10 replications. I compare five methods: ordinary least squares (OLS), Random Forest with 500 trees and $\lfloor P/3\rfloor$ variables per split, a Multi-Layer Perceptron (MLP) with three hidden layers of 200 units each, ReLU activations, dropout regularization (rate =0.2=0.2), trained via Adam with early stopping (patience =20=20 epochs, 15% validation holdout), Gradient Boosting with 500 trees and learning rate 0.01, and Attention Regression with five heads and regularization $\lambda=10^{-3}$. Performance is measured by out-of-sample $R^{2}$, which is informative because it is comparable across DGPs and can be analyzed in relation to the maximal theoretically attainable $R^{2}$, which is $R^{2}_{\max}=\text{SNR}/(\text{SNR}+1)$.
Figure 2: Model Performance Comparison Across Data Generating Processes
Notes: This figure displays the average $R^{2}$ values for each model across all experimental conditions (sample sizes $N\in\{500,1000,2500,5000\}$ and signal-to-noise ratios $\text{SNR}\in\{0.5,1.0,2.0,3.0\}$). Error bars represent $\pm 1$ standard deviation across the 16 conditions within each DGP. The “Overall” category aggregates performance across all six DGPs (96 total conditions). Values above bars indicate the mean $R^{2}$ of the best-performing model for each DGP. The dashed vertical line separates the aggregate results from individual DGP performance.
Results.
As summarized in Figure 2 and detailed in Table LABEL:tab:benchmark_results, the MLP achieves the highest overall $R^{2}$ (0.52) across all 96 experimental conditions, followed by Attention Regression (0.48), Random Forest (0.45), Gradient Boosting (0.42), and OLS (0.28). The MLP’s strong performance is expected: as a universal function approximator with flexible hidden representations, it can adapt to arbitrary nonlinearities given sufficient data.
What is more notable is how competitive Attention Regression remains despite its simpler architecture. It consistently outperforms both Random Forest and Gradient Boosting overall and in four of the six DGPs. On Friedman1, it wins outright (0.53 vs. 0.52 for MLP and 0.49 for RF/GBM), and it essentially ties the MLP on Friedman3 (0.54 for both) while beating RF and GBM by 4 and 1 percentage points respectively. On Friedman2, Attention Regression outperforms Random Forest by 7 percentage points and Gradient Boosting by 15 points, though the MLP leads by about 5 points. These Friedman designs rely heavily on interactions and smooth nonlinear transformations—exactly the patterns that learned attention weights can capture through adaptive similarity weighting.
The expected exception is the Linear DGP, where OLS dominates (0.56) due to correct specification. Here, all flexible methods pay a price for their flexibility, with Attention Regression trailing by about 8 points. On the periodic and radial DGPs (Rotated Sine and Soft Radial), Attention Regression beats both tree-based methods on Rotated Sine (0.41 vs. 0.38 for RF and 0.25 for GBM) while the MLP’s advantage is more pronounced, reflecting its ability to approximate high-frequency oscillations through compositions of ReLU activations. On Soft Radial, Random Forest leads among non-neural methods (0.47), with Attention Regression (0.46) close behind.
As visible from the error bars in Figure 2, variability is higher for Attention Regression and MLP in small samples or low SNR settings, but this gap narrows with larger $N$, consistent with typical sample-efficiency patterns of flexible nonlinear models.
These simulations demonstrate that a straightforward implementation of Attention Regression performs remarkably well relative to standard nonlinear ML benchmarks. All models were run with basic, off-the-shelf hyperparameters—the purpose of the paper is not to introduce yet another nonlinear estimator, but to clarify the connection between attention mechanisms and least squares. Still, it is notable that taking the formulation literally and estimating $\boldsymbol{\Omega}$ directly produces a model that often matches tree-based methods and remains competitive with neural networks.
6 Discussion and Directions for Future Research
I show that ordinary least squares predictions can be reformulated as the output of a constrained attention module, similar to those underpinning modern large language models. This perspective extends the interpretation of attention beyond its standard information retrieval framework, bridging it with classical statistical methods and making it more intuitive for researchers familiar with traditional regression analysis.This is facilitated by reframing OLS as a similarity-based method in a transformed orthonormal space, where the solution to the optimization problem also coincides with that of an optimal embedding problem. I then discuss how key features of transformer architectures—such as dimensionality reduction, nonlinearity, multi-head attention, self-attention, and masking—can be understood through the lens of OLS.
The "OLS perspective" suggests that attention mechanisms can be understood as a nonlinear least squares method with generated features, where learned concepts replace raw predictors. These concepts emerge from a dimension-reduced representation of the input, akin to transformed tokens or feature embeddings. Instead of forming direct mappings, the attention mechanism restructures prediction as a correlation-based problem in this learned space, much like any nonparametric supervised learning approach. [^2] This view naturally aligns with the rest of the deep learning canon, where learned latent or hidden representations are a hallmark of the model’s ability to capture complex patterns. In certain contexts, these latent concepts can be extracted and visualized to enhance interpretability (kim2018interpretability; goulet2024neural).
The output of an attention module, often termed a contextualized embedding, integrates information across a sequence through projection rather than explicit sequential modeling. This aligns with time series econometrics tools like VARs, where equations can be estimated separately via OLS as if they were cross-sectional data, once lags are encoded. Each token’s contextualized embedding can be seen as a fitted value in a model where tokens simultaneously "explain" and are "explained by" others.
A macroeconomic parallel is the interpretation of high interest rates across inflationary environments: their meaning shifts depending on context, much like attention-based embeddings adjust dynamically. Context for interest rates is often formalized through a Taylor rule, where higher inflation and stronger real activity justify higher rates. Rates that align with the rule carry different implications than those that deviate significantly—known as monetary policy shocks. Again, "context" takes the form of expected conditions at $t$ given available information.
Word Shocks?
Macroeconomists are particularly fond of shocks, the difference between the realized value of $y_{t}$ and its expected value. In contrast, LLMs construct conditional means as indicators of context, serving as a form of feature engineering for subsequent feed-forward neural network modules. In this spirit, it is intriguing to consider the possibility of word shocks. One could examine the difference between the original, context-independent embeddings and their context-conditioned versions. This difference could be interpreted as a "shock" or a granular measure of linguistic surprise–capturing unexpected word choices or unusual word adjacencies. Some factor model or supervised aggregation scheme (as in, e.g., albacore) could be leveraged to map this back into macroeconomic series of interest.
This granular approach would differ significantly from methods that construct a single shock series by regressing an observed time series (such as short-term interest rates) on sentiment extracted from text (aruoba2024identifying). Instead, each word in a sentence could be decomposed into the sum of its contextual contribution and a shock component. Intermediate aggregation schemes could also prove useful, such as ranking policy reports on a spectrum from "utterly predictable" to "jarring." Some existing work moves in this direction, such as fischer2023fed, who tracks various aspects of Fed communications, including their opacity level.
Proximity-based Identification Schemes?
From the OLS sides of things, its view as a linear attention module, where similarity scores are derived from proximity in an orthonormal space, also suggests interesting research questions. For instance, in econometrics and statistics more broadly, the estimator $\boldsymbol{\beta}$ is considered unbiased under the standard exogeneity condition, $\mathbb{E}[\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{\varepsilon}]=0.$ where $\boldsymbol{\varepsilon}\in\mathbb{R}^{N}$ is the vector of true errors in the data-generating process. Given our alternative framing of OLS as an attention mechanism, it is natural to ask whether this perspective can serve as a jumping board for alternative identification restrictions. Specifically, can we define looser conditions through $\boldsymbol{\Omega}$ under which unbiased estimation is still achievable, even when standard exogeneity does not strictly hold? This question is particularly relevant since, although this paper has focused on predictions, the coefficient vector $\hat{\boldsymbol{\beta}}$ itself can be expressed as the difference between two predictions.
Speaking in terms of similarity naturally draws a parallel to matching estimators in causal inference. Although these estimators — whether based on nearest neighbors, kernels, or propensity scores — are constructed through similarity principles, their validity ultimately relies on the unconfoundedness assumption. This condition, a close cousin of the usual exogeneity assumption, is grounded in a correlation-based view of the problem rather than purely in terms of similarity. The connection between OLS estimation of $\hat{\boldsymbol{\beta}}$ and proximity-based causal estimators is further explored in dual2.
Recent Developments.
Since the initial circulation of this paper, the landscape of large language models has continued to evolve rapidly. Reasoning-focused architectures and retrieval-augmented generation (RAG) systems have gained prominence, yet the attention mechanism remains their computational backbone. Alternative architectures such as Mamba (gu2023mamba) and FFT-based methods (fein2025fft) have emerged as potential successors, but attention-based Transformers continue to dominate deployed systems. The OLS perspective developed here remains applicable to these newer variants insofar as they rely on similarity-based weighting of stored representations or can be framed as adaptive filters in a latent space.
Limitations and Closing Remarks.
The analogy proposed in this paper itself requires contextualization. Some features of the Transformer architecture were not formally discussed. For instance, the output of the attention module is not directly compared with a raw supervision target. Instead, the target is also encoded, and the model learns to align these representations through further layers. Moreover, residual connections and layer normalization play crucial roles in stabilizing training, preserving gradient flow, and maintaining consistent representations across layers.
Thus, it is important to clarify—at the elevated risk of stating the obvious—that this paper does not assert that LLMs are merely performing ordinary least squares. What it asserts, however, is that the principles underlying the exceptional capabilities of Transformer architectures, particularly through their attention modules, are aligned with least squares principles and, as such, are not as distant from the traditional statistical toolkit as they might initially appear. Hopefully, this understanding can prove useful to both lines of work.
Appendix A Detailed Simulation Results
Table 1: Model Performance Comparison (Out-of-Sample $R^{2}$) Across Data Generating Processes
Models
DGP NN SNR OLS RF MLP GBM Att Reg
[…]
1000 0.5 0.32 0.27 0.31 0.28 0.29
1 0.49 0.44 0.47 0.45 0.42
2 0.66 0.60 0.64 0.62 0.55
3 0.74 0.68 0.73 0.71 0.68
[…]
5000 0.5 0.34 0.31 0.33 0.32 0.26
1 0.50 0.47 0.49 0.48 0.38
2 0.67 0.64 0.66 0.65 0.52
3 0.75 0.72 0.75 0.73 0.57
[…]
1000 0.5 0.25 0.27 0.29 0.27 0.30
1 0.37 0.42 0.45 0.43 0.47
2 0.50 0.58 0.62 0.59 0.64
3 0.56 0.66 0.71 0.67 0.73
[…]
5000 0.5 0.25 0.30 0.32 0.30 0.33
1 0.37 0.46 0.48 0.45 0.48
2 0.49 0.62 0.65 0.61 0.64
3 0.55 0.70 0.73 0.68 0.72
[…]
1000 0.5 0.02 0.18 0.27 0.12 0.23
1 0.04 0.30 0.44 0.23 0.41
2 0.05 0.42 0.61 0.34 0.57
3 0.06 0.49 0.70 0.40 0.63
[…]
5000 0.5 0.04 0.26 0.31 0.20 0.27
1 0.05 0.41 0.47 0.30 0.41
2 0.07 0.55 0.64 0.41 0.52
3 0.08 0.63 0.73 0.46 0.63
[…]
1000 0.5 0.30 0.27 0.30 0.29 0.31
1 0.45 0.43 0.47 0.46 0.48
2 0.60 0.59 0.64 0.63 0.65
3 0.67 0.67 0.73 0.71 0.74
[…]
5000 0.5 0.29 0.29 0.31 0.31 0.31
1 0.44 0.46 0.48 0.47 0.48
2 0.59 0.63 0.65 0.64 0.65
3 0.66 0.71 0.73 0.72 0.74
[…]
1000 0.5 0.10 0.18 0.27 0.12 0.21
1 0.15 0.30 0.43 0.20 0.37
2 0.20 0.42 0.61 0.30 0.51
3 0.23 0.48 0.70 0.35 0.56
[…]
5000 0.5 0.10 0.26 0.30 0.16 0.22
1 0.16 0.40 0.47 0.26 0.36
2 0.21 0.54 0.64 0.35 0.47
3 0.24 0.61 0.73 0.40 0.58
[…]
1000 0.5 -0.00 0.26 0.27 0.23 0.21
1 -0.00 0.40 0.45 0.36 0.38
2 -0.00 0.55 0.62 0.49 0.55
3 -0.00 0.62 0.70 0.56 0.63
[…]
5000 0.5 -0.00 0.29 0.31 0.25 0.29
1 -0.00 0.45 0.48 0.39 0.45
2 -0.00 0.61 0.65 0.52 0.56
3 -0.00 0.69 0.73 0.59 0.62
Notes: Each cell reports out-of-sample $R^{2}$. Models: OLS (Ordinary Least Squares), RF (Random Forest), MLP (Multi-Layer Perceptron), GBM (Gradient Boosting), Att Reg (Attention Regression). Green bold = best; black bold = second-best.